OpenStudy (anonymous):

If f(x) = x2 + x and g(x) = x − 3, find f(g(7)). A)20 B)28 C)32 D)53

OpenStudy (anonymous):

@Ghostedly

OpenStudy (ghostedly):

It's 20.

OpenStudy (anonymous):

explain how

OpenStudy (ghostedly):

First, compute g(7)

OpenStudy (ghostedly):

To do that, plug x=7 into g(x)=x-3

OpenStudy (ghostedly):

g(7)=7-3=4

OpenStudy (ghostedly):

So, g(7)=4

OpenStudy (ghostedly):

Then you plug g(7) into f(x) to get f(g(7))=f(4)=^2+4=16+4=20

OpenStudy (ghostedly):

So f(g(7))=20

OpenStudy (ghostedly):

@desireesmith15

OpenStudy (anonymous):

huh

OpenStudy (ghostedly):

There's the explanation.

OpenStudy (anonymous):

thanks i got it i medal u too